OpenStudy (anonymous):

The grades on the last math exam had a mean of 88%. Assume the population of grades on math exams is known to be normally distributed with a standard deviation of 6. What percent of students earn a score between 72% and 90%? (5 points) 0.6255 0.5843 0.6754 0.5346

OpenStudy (anonymous):

explain also how to do it

OpenStudy (kropot72):

This can be done using a standard normal distribution table. First you need to find the z-scores for 72 and 90. Can you do that?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

I want to know how to do this, the teacher just gives us homework and tells us to figure it out by ourselves.

OpenStudy (anonymous):

@wio

OpenStudy (kropot72):

You can use the standard normal distribution table here if you need it http://lilt.ilstu.edu/dasacke/eco148/ztable.htm The z-scores are found using the formula: \[z=\frac{X-\mu}{\sigma}\] The z-score for 90 is: \[\frac{90-88}{6}=0.33\] And the z-score for 72 is: \[\frac{72-88}{6}=-2.67\] Now you need to use a standard normal distribution table to find the cumulative probabilities for z-scores of 0.33 and -2.67. Please do that and post your results.

OpenStudy (anonymous):

Is the answer C?

OpenStudy (kropot72):

You need to find the two values that I asked for. Then we can quickly find the solution.

OpenStudy (anonymous):

.0038 and .6293

OpenStudy (kropot72):

Good work! Now just do this calculation to find the percent of students who earn a score between 72% and 90%: \[(0.6293-0.0038) \times100=\ ?\ percent\]

OpenStudy (anonymous):

A

OpenStudy (kropot72):

Actually the question asks for a percentage, but the answer choices are decimal fractions. Yes, A is correct :)